Math 301 — Introduction to Probability and Statistics I, Fall 1996

Professor: Bret Larget Office: 432 College Hall Phone: 396-6469 (Office)

396-6467 (Department)

Office Hours: Tuesday 3:00 – 5:00 P.M.

Thursday 1:30 – 2:30 P.M. or by appointment

Informal Office Hour: Monday 12:00 – 1:00 P.M. in the Options Cafeteria Class Hours: Tuesday and Thursday, 11:30 A.M. – 12:45 P.M. Room: 225 College Hall Text: Mathematical Statistics and Data Analysis 2^{nd} Edition by John Rice

Prerequisites: You must have passed Math 116 (or it’s equivalent) for this course. Integration and differentiation are used heavily.

Homework (20%): Several assignments consisting of textbook problems and a few assignments using the computer will be given during the semester. Part of this grade will include demonstrating homework solutions in class. You are permitted and, in fact, encouraged to work with other students on your homework assignments. However, each student must turn in their own assignment.

Midterm (20%): There will be one 75-minute closed-book midterm examination on October 8. Part of the midterm will be a take-home examination.

Quizzes (15%): There will be five 20-minute quizzes. Dates will be announced at least one week prior to the quiz.

Class Project (15%): A substantial project including a three to ten page written summary. Project ideas will be discussed in class.

Final (30%): The comprehensive closed-book final examination will be on Friday, December 13 from 1:15 to 3:15 P.M.

Grading: Final grades will be based on the percentages above. Final grades will not be given according to a fixed scale. Letter equivalents will be announced with each test and quiz. The +/- system will be used.

Important Dates:

Tuesday, October 8 Midterm Examination

Tuesday, December 3 Project Reports

Thursday, December 5 Project Reports, Last Day of Class Tuesday, December 10 Last day written project may be turned in Friday, December 13 Final examination

Academic Honesty: A student found cheating on a test or assisting others in the class on a test will receive an F for the course and may face further sanctions. Copying another student’s assignment is considered cheating and will result in receiving a 0 for that assignment.

Math 301 — Course Outline, Fall 1996 (Subject to Change)

Date  Topics  Sections to Read  Important Events

Aug 27 Random Variables, Distributions, Outcome Spaces 1.1 – 1.3, 2.1 – 2.2 Aug 29 pmf’s, pdf’s, and cdf’s

Sep 3  Coin Tossing Distributions  2.1.1 – 2.1.3
Sep 5  Counting Problems  1.4  Quiz

Sep 10 Hypergeometric Distribution / Sampling 2.1.4 Sep 12 Poisson Distribution 2.1.5 Sep 17 Conditional Probability / Independence 1.5 – 1.6 Sep 19 Continuous Random Variables 2.2.1 – 2.2.3 Quiz Sep 24 Poisson Process 2.1.5 Sep 26 Functions of Random Variables 2.3 Oct 1 Discrete Joint Distributions 3.1 – 3.2 Oct 3 Continuous Joint Distributions 3.3 Oct 8 Inclass Midterm Oct 10 Conditional Distributions 3.5 Take-home Midterm Oct 15 Functions of Joint Distributions 3.6 Oct 17 Order Statistics 3.7 Oct 22 Expectation 4.1 Oct 24 Expectation of Functions of Random Variables 4.1.1 – 4.1.2 Quiz Oct 29 Variance and Standard Deviation 4.2 Oct 31 Covariance and Correlation 4.3 Nov 5 Law of Large Numbers and CLT 5.1 – 5.3 Quiz Nov 7 Distributions Derived from the Normal 6.1 – 6.3 Nov 12 Simple Random Sampling 7.1 – 7.3 Nov 14 Normal Approximation 7.3.3 Nov 19 Catch-up Day Nov 21 Review Quiz Dec 3 Project Reports Dec 5 Project Reports

Final Examination on Friday, December 13